# SPECIAL RELATIVITY

### v9

@@ 9.5 Metrics in the Accelerated Frame
Let us follow out the consequences of assuming that the inertial frame has no accelerational history and the (observer) space-time metric at rest has diagonal form. We start with the two emitters in the inertial frame. The first emitter is at rest and the second is moving with velocity $v$. As a result of Galilean boost, we obtained the metric of the accelerated emitter in the inertial frame Eq. (11). In this coordinatization, we describe the non-accelerated light source using diagonal metric, $ds^{2}=c^{2}dt^{2}-dx^{2}$.
@@
Let us now return to our consideration of the accelerated frame. In the accelerated frame we have the reciprocal velocity. Second emitter is at rest and the first is moving with velocity $-v$. Now what about reciprocity in physics? According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated frame and the inertial frame and the accelerated observer should obtain diagonal metric for the second emitter at rest in the accelerated frame. The principle of reciprocity implies that the metric Eq. (15) describes the aberration of light effect in the case of the first (moving) emitter in the accelerated frame. The metric of the moving emitter in the accelerated frame is the same as the metric Eq. (11) of the moving emitter in the inertial frame. According to principle of reciprocity, the difference is only that the sign of velocity $v$. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of electrodynamics equations does not dictate the Lorentz covariance of the solutions.
@@
The commonly accepted approach to special relativity does not account for the inertial (pseudo-gravitational) force within the accelerated frame as the agency of asymmetry. Space-time geometrical analysis of the accelerated frame (see Section 9.1) leads to the asymmetry (non-reciprocity) between the inertial frame and the accelerated frame. We obtain in the accelerated frame the Langevin metric Eq. (15) of the second emitter. This metric describe the electrodynamics equations in the accelerated frame. We found of the essential asymmetry between inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Because we are using absolute time coordinatization, the initial conditions (i.e. radiation wavefront orientations) are identical in both frames.
We are going to discuss the asymmetry between inertial and accelerated frames. In the discussion up to this point, we assumed absolute time coordinatization. The Langevin metric, together with initial condition, describes the aberration of light effect for the second emitter at rest in the accelerated frame. The essential asymmetry is rooted in the electrodynamics equations. However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame. The explanation consists in using a metric diagonalization procedure. The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric for the second emitter at rest and the (Langevin) metric Eq. (15) for the first (moving) emitter. Now we see that there is a symmetry (reciprocity) in electrodynamics equations between the accelerated and the inertial frame. This is a consequence of the pseudo-Euclidean geometry of space-time. Our impossibility to detect orbital velocity using second order optical (interference) experiments is a result of this metric reciprocity, see Chapter 11 for more details. Now the question is, where asymmetry comes from? Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions. The plane radiation wavefront of both emitters rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization.
The previous derivation of the metrics in the accelerated frame includes one delicate point. According to a reciprocity in metrics, time dilation is symmetric. This contradiction is resolved by noting that metric diagonalization procedure does not preserve continuity of the metric tensor. In the previous chapters we learned that the accelerated frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one. In fact, the Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. We shall discuss the smooth tailoring of the metric tensor further in the next section.
9.6 Mathematical Expansion of Special Relativity onto Acceleration
Before we begin the next topic, we would like to describe a number of mathematical ideas that are used in the literature of time dilation effect. Suppose we have two clocks in the inertial frame $S$ without accelerational history. Let first of these clocks always be at rest in the frame $S$ and the second clock accelerates and start to move with constant velocity $v$ along the $x$ axis. The proper time of the first clock coincides with the coordinate time $d\tau=dt$. Since the motion of both clocks is uniform, frames $S$ and $S_{n}$ will be inertial. According to textbooks, the coordinates and time in this frames should be related by the Lorentz transformation and the metrics of both frames will be diagonal. At first site the problem is complete symmetrical and there is no difference in the metrics in the frames $S$ and $S_{n}$. The situation seems paradoxical.
We know that independent of the chosen coordinatization the proper time of the second accelerated clock is always less than that of the first clock. The time dilation of the accelerated clock is an absolute effect but not relative. It is independent of the time coordinate definition at the inertial part of the trajectory. This is an intrinsic property of space-time geometry. A more careful analysis shows that one must take into account that the metric tensor must be a continuous quantity.
In the previous chapters we demonstrated that many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization (i.e non diagonal metrics). We will keep the mathematical complexity to a minimum and assume that the second clock begins at $t=0$ to move with constant acceleration $a>0$ and does so till $t=T$. Now we would like to discuss an expression for the velocity from the viewpoint of inertial (non accelerated) observer. We know that uniformly accelerated motion must obey the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$. This is readily integrated $\gamma\vec{v}=\vec{a}t$. We will then have $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$. We emphasize that $dx(t)/dt$ is the velocity of the second clock as seen by the inertial observer. Under the inverse Galilean transformation $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$, $dt_{n}=dt$ the metric of the uniformly accelerating observer becomes (3)
At $t=0$ this metric naturally coincides with the $S$ frame metric and we have $g_{00}=1$. But here at the inertial part of the trajectory the metric of both frames will be diagonal in the Lorentz coordinatization and so at $t=T$ the metric tensor of the accelerated frame must abruptly change from the value $g_{00}=(1+a^{2}t_{n}^{2}/c^{2})^{-1}$ to the value $g_{00}=1$. This shows that we must match the coordinates in the accelerational and inertial parts of the (second clock) trajectory.
During acceleration we picked the absolute time coordinatization. Therefore, for the metric tensor to be continuous when the reference frame passes from non-inertial motion to inertial one we will make use of the metric Eq. 19. We will therefore require that at $t=T$ the metric of a non inertial frame associated with the second clock would pass continuously into the Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},t_{n})$ to the coordinates $(x,t)$ of the inertial observer by the Galilean boost $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$. In this case, in the interval of coordinate time $T<t_{n}$ the reference frame of the second clock has the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$. We emphasize that $v$ is the velocity of the second clock as seen by the inertial observer at the second (inertial) stage of motion. It becomes obvious that at $t=t_{n}=T$ the metric of non inertial frame goes into the metric of the (accelerated) inertial frame in a continuous manner.
9.7 Bibliography and Notes
1. For a general discussion of the equivalence principle we suggest reading the paper [42].
2. In the general case, the problem to assigning Lorentz coordinates in an inertial frame is complicated. Let us consider a source, arbitrary accelerating in the inertial frame, and let us analyze its evolution with a Lorentz coordinate system. The permanent rest frame of the source is obviously non inertial. To get around that difficulty, one introduces an infinite sequence of comoving coordinate systems. At each instant, the comoving coordinate system is a Lorentz coordinate system centered on the source and moving with it. As the source velocity changes to its new value at an infinitesimally latter instant, a new Lorentz coordinate system centered on the source and moving with it at the new velocity is used to observe the source. We should make one further remark about this covariant algorithm. An opinion is sometimes expressed that described above algorithm includes a hidden postulate. It seems necessary a dynamical assumption to justify attributing to an accelerated clock the same rate as a clock in inertial motion in relation to which it is momentary at rest. This is, in view of some authors, an extra condition that a clock must satisfy. It is assumed that the effect of the motion on the clock depends only on it instantaneous speed, not its acceleration. This condition often refereed to as a ”clock hypothesis”. To quote Brown [17]: ”If the accelerating forces are small in relation to the internal restorative forces of the clock, then the clock’s proper time will be proportional to the Minkowski distance along its world line. … This condition is often referred to as the clock hypothesis, and its justification, as we have seen, rests on accelerative forces being small in the appropriate sense.” We state that the clock hypothesis does not have status of independent hypothesis is not needed as an independent postulate in the theory of relativity. As discussed above, the transformations within a single inertial frame are simply another parametrization of the observations of the inertial observer. According to such transformations, the observer (and his clocks) motion relative to the fixed stars does not changes. In other words, when we discuss only the observations of an inertial observer the problem of accelerated clock does not exist at all.
3. For a general discussion of the mathematical expansion of special relativity onto acceleration we suggest reading the book [16].
10 Relativistic Measurement
In our approach to physics in general and of the theory of relativity in particular we think it very important always to remember that we are dealing with objective physical quantities measured by the measurements. We illustrate the problem of measurement in the special relativity by typical examples.
@@ 10.1 The Notion of Time in Special Relativity Theory
There are many changes in what concepts are important when we go from classical to relativistic theory. We have already discussed some of them in previous chapters. We will show the connection between the classical notion of time and the relativistic time. The theory of relativity show us that the notion of time is not what we would have expected on the basis of our intuitive ideas.
@@
We attempt to give a definition of time on the basis of a principle ”Each physical magnitude is defined by the method it is measured.” Let us see how this definition works. In the case of length we know that there exist a length standard and we know that any body has its particular length. It should be note, however, that we are dealing with objective physical length only in the case of a stationary object. In contrast to this, the length of a moving object (e.g. moving rod) is convention-dependent (i.e. synchronization-dependent) and has no exact objective meaning. Now we are ready to consider the time problem in special relativity theory.
@@
Here is one example which shows in a circumstance that is easy to understand. Suppose we know the law of muon decay in the Lorentz rest frame. When a Lorentz transformation of the decay law is tried, one obtains the prediction that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$. We can interpreted this result by saying that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would reduced to 1/2 of the origin population. The time measured in this case is ”length” which is metrized as described above. That leads us to interesting question. Is the same statement perhaps also true for all cases of time measurements in the framework of special relativity? We claim that the answer to this question is positive.
@@
Our next example has to do with frequency. What about the frequency measurements? Let us suppose that we have a (Fabry-Perot) interferometer. The frequency measured in this case is (standing wave) ”length”. Frequency measurements automatically obey the relationship between the frequency and interference. Another example of the relation between interference and frequency measurements is a grating spectrometer. The frequency measurement in this case is a position of the light spot along the dispersion direction. There is no specific, important physical difference between diffraction grating and interferometer. Grating spectrometer is also based on the interference of incident and reflected light waves.
@@ 10.2 Invariance of the Two-Way Speed of Light and the Interference Phenomena
We must now discuss a certain feature of the phenomenon of interference. Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes. One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection.
@@
Let us now consider a reflection and refraction in the context of special relativity. We are now in a position to understand what is happening when light radiated by a moving emitter goes through a stationary material like glass. We see that what we have to do is to calculate the interference between the incident radiation and the radiation due to the material. At first glance, if one wants to calculate the refraction and the reflection in a single inertial frame one should take into account that the metrics, and consequently the electrodynamics equations, are different for a moving emitter and a stationary glass. So the possibility of calculating interference effects is not clear.
@@
Figure 47: Light radiated by a moving emitter reflects from a stationary mirror in the collinear geometry. There is interference between the incident radiation and the radiation due to the material. Galilean transformation of the velocity of light is consistent with the electron-theoretical explanation of refraction and reflection. The Minkowski metric and the metric Eq. (11) give the same prediction for the wavelength of the standing wave.
@@
If the mirror is at rest and the emitter is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the same metric should be applied and kept in a consistent way for both incident and scattered waves. The requirement that the electrodynamics equations must be identical for both incident and reflected waves appear to be a paradox here. In fact, we have two different metrics.
@@
We have already made a few remarks about two different metrics in the aberration of light problem. In the Chapter 4, we considered only the first order approximation. We pointed out that the paradox is avoided because the peculiarity of the aberration of light (perpendicular) geometry is that even after the Galilean transformation the electrodynamics equations in the absolute time coordinatization will have Maxwell’s form for the incident plane wave. Let us now see what happens in the collinear geometry, Fig. 47.
@@
Here we just summarize the general idea. We will show that the paradox is avoided because Minkowski metric and metric Eq. (11) give the same prediction. The point is that all methods to measure the interference, indeed, the standing wave (i.e. round-trip) measurements. In contrast, the deviation of the energy transport direction is a geometrical effect. Because we have empirical access only to the round-trip average speed of light the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning.
@@
We must now discuss a certain feature of the phenomenon of interference. One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection. It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different. This is simply not true in physics. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase which varies with position in space. This affects the velocity and amplitude envelope of the single wave which results from the superposition of the two separate waves.
@@
The following simple analysis confirms these ideas. We will describe effect by assuming that the metric is non-diagonal, Eq. 11. Let $\exp i(\omega t-kx)$ represent an incoming wave whose velocity is $\omega/k=c+v$. Similarly, let $\exp i(\omega t+k^{\prime}x+\phi)$ represent another out-coming (scattered radiation) wave of the same amplitude and the same frequency, a different velocity $\omega/k^{\prime}=c-v$ and different phase. The superposition of these two waves is represented by $\exp i(\omega t-kx)+\exp i(\omega t+k^{\prime}x+\phi)=2[\cos[(k+k^{\prime})x/2+\phi/2]]\exp i[\omega t-(k-k^{\prime})x/2+\phi/2]$. There is the cosine factor representing an amplitude envelope which is stationary in space and whose periodicity is inversely proportional to the difference in the propagating constants $k$ and $k^{\prime}$ of the two component waves. This can be written in a simpler form $2[\cos[\omega x/[c(1-v^{2}/c^{2})]+\phi/2]]\exp i[\omega t-xv\omega/[c^{2}(1-v^{2}/c^{2})]+\phi/2]$.
@@
Suppose that the source at rest is emitting waves at frequency $\omega_{0}$. In the lab frame after the Galilean transformation the velocity of incoming wave is $c+v$. Thus if $\omega_{0}$ is the natural frequency, the observed in the lab frame frequency would be $\omega=\omega_{0}/[1-v/(c+v)]$. Therefore, the observed in the lab frame frequency is $\omega=\omega_{0}(1+v/c)$. The shift in frequency observed in the above situation is the well known Doppler effect. Our equation for superposition of two waves now looks like $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
@@
Suppose that an observer in the laboratory performs the standing wave measurement. We should examine what parts of the measured data depends on the choice of synchronization convention and what parts do not. We state that time oscillation has no intrinsic meaning - its meaning only being assigned by a convention. In particular, one can see the connection between the time shift $xv\omega_{0}(1+v/c)/c^{2}$ in $\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$ and the issue of distant clock synchrony. Note that the scale of time (frequency) is also unrecognizable from physical viewpoint.
@@
Suppose that the laboratory observer performs a measurement of the wavelength of the standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ gives the connection between the spatial line element and the metric. With the help of Eq.(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the inertial frame. This is explained as due to Lorentz contraction of the measuring rods in the inertial frame. Then, when the measured data is analyzed, the laboratory observer finds that the wave number is equal to $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/\sqrt{1-v/c}$. The observer finds that the wavelength of radiation from moving source (the source moves towards the observer) is decreased by the factor $\sqrt{1-v/c}/\sqrt{1+v/c}$. We see that it is the same factor that we can obtain by assuming that the metric is diagonal. In other words, the laboratory observer will measure the same two-way speed of light, irrespective of the choice of metric.
@@
The theory of relativity shows us that there is a certain quantity, which we call two-way speed of light, that does not change under any reference frame transformations. Another way in which our complicated problem can be solved is the following interesting one. The reason why the laboratory observer will measure the same two-way speed of light is that the method is based on the phenomenon of interference. There is a fact governing all interference phenomena. The law is called the relativistic invariance of phase. It states that there is a certain quantity, which we call phase, that does not change under any reference frame transformations. The relativistic invariance of phase is enormously useful in making analysis (as we will illustrate in a few examples in the Chapter 11) without knowing all the formulas.

### v10

@@ 12.5 Metrics in the Accelerated Frame
Let us follow the consequences of assuming that the inertial frame has no accelerational history and the (observer) space-time metric at rest has a diagonal form. We start with the two emitters in the inertial frame. The first emitter is at rest and the second is moving with velocity $v$. As a result of the Galilean boost, we obtained the metric of the accelerated emitter in the inertial frame Eq. (11). In this coordinatization, we describe the non-accelerated light source using diagonal metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
@@
Let us now return to our consideration of the accelerated frame. In the accelerated frame we have the reciprocal velocity. The second emitter is at rest and the first is moving with velocity $-v$. Now what about reciprocity in physics? According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated frame and the inertial frame and the accelerated observer should obtain a diagonal metric for the second emitter at rest in the accelerated frame. The principle of reciprocity implies that the metric Eq. (15) describes the aberration of light effect in the case of the first (moving) emitter in the accelerated frame. The metric of the moving emitter in the accelerated frame is the same as the metric Eq. (11) of the moving emitter in the inertial frame. According to the principle of reciprocity, the difference is only that the sign of velocity $v$. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
@@
We are going to discuss the asymmetry between inertial and accelerated frames. Space-time geometrical analysis of the accelerated frame leads to the asymmetry (non-reciprocity) between the inertial frame and the accelerated frame. We obtain in the accelerated frame the Langevin metric Eq. (15) of the second emitter. This metric describes the electrodynamics equations in the accelerated frame. We found the essential asymmetry between inertial and accelerated frames, namely, Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Because we are using absolute time coordinatization, the initial conditions (i.e. radiation wavefront orientations) are identical in both frames  [^53] Many physicists tend to think that the validity of Maxwell’s laws in all inertial frames implies the elimination of the privileged frame. To quote e.g. Dieks [51]: ” As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”.
We would now like to describe an apparent paradox. The argument that all inertial frames are equivalent runs something like this. We know that to keep Lorentz coordinates in the accelerated frame we need to perform a metric diagonalization. Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in these frames should be related by the Lorentz transformation. At first sight, the problem is completely symmetrical and both frames are equivalent. However, there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. Indeed, we should expect the metric tensor to be changed. The difference between the inertial and accelerated frames is understandable. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$. We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve the continuity of the metric tensor we have no symmetry between inertial frames.
13 Relativistic Measurements
@@ 13.1 The Notion of Time in Special Relativity Theory
The theory of relativity shows us that the notion of time is not what we would have expected on the basis of our intuitive ideas. We attempt to give a definition of time on the basis of a principle ”Each physical magnitude is defined by the method it is measured.” Let us see how this definition works. In the case of length, we know that there exists a length standard and we know that anybody has a particular length. It should be noted, however, that we are dealing with objective physical length only in the case of a stationary object. In contrast to this, the length of a moving object (e.g. moving rod) is convention-dependent (i.e. synchronization-dependent) and has no exact objective meaning. Now we are ready to consider the time problem in special relativity theory.
@@
Here is one example which shows a circumstance that is easy to understand. Suppose we know the law of muon decay in the Lorentz rest frame. When a Lorentz transformation of the decay law is tried, one obtains the prediction that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$. We can interpret this result by saying that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would reduced to 1/2 of the origin population. The time measured in this case is ”length” which is metrized as described above. That leads us to an interesting question. Is the same statement perhaps also true for all cases of time measurements in the framework of special relativity? We claim that the answer to this question is positive.
@@
Our next example has to do with frequency. What about the frequency measurements? Let us suppose that we have a (Fabry-Perot) interferometer. The frequency measured in this case is (standing wave) ”length”. Frequency measurements automatically obey the relationship between frequency and interference. Another example of the relation between interference and frequency measurements is a grating spectrometer. The frequency measurement in this case is the position of the light spot along the dispersion direction. There is no specific, important physical difference between diffraction grating and interferometer. The grating spectrometer is also based on the interference of incident and reflected light waves.
Now let us see how we can compare the clock rates in the initial inertial frame $S$ and accelerated frame $S_{n}$. It is generally believed that observing clock rates in the two frames $S$ and $S_{n}$ moving relative to each other, one can only compare readings of one clock from one frame with readings of several clocks from another frame because two clocks from different frames get together at the same point in space only once. In one of the frames, there must be at least two clocks that are supposed to be synchronized.
We have already pointed out that in all cases of time measurements in the framework of special relativity the time measured in this case is actually ”length”. In the case of length measurement of resting objects, we are dealing with objective (i.e. convention independent) physical quantity. Because we have empirical access only to the length measurements, it is impossible to agree with the textbook statement that one can only compare readings of one clock from one frame with readings of several synchronized clocks from another frame. Let us try out our operational interpretation of time measurements on the light clock example, to see how it works.
Let us consider a specific thought experiment. We shall discuss the situation where time marks can be imprinted on the moving object (screen) by a clock. First, we describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock. A screen is at rest in the initial inertial frame. Suppose that the top mirror is semitransparent for marking the (moving) screen to measure the time between clock clicks. In the case of Langevin metric Eq.(15), the speed of light radiated by the accelerated emitter in the transverse direction is $c\sqrt{1-v^{2}/c^{2}}$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. The spacing of the time marks, i.e the distance $\Delta x_{n}$ through which the object (screen in the initial inertial frame) falls in the time is $\Delta x_{n}=\Delta x=2lv/[c\sqrt{1-v^{2}/c^{2}}]$.
Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=dt=0$ that $dx_{n}=dx$, i.e. the length of the physical rod in the inertial frame coincidence with the coordinate distance in the accelerated frame. Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the initial inertial frame, this implies that, as physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame. The accelerated clock rate measured by the inertial observer is actually spacing of the time marks on the screen in the initial inertial frame.
Second, we wise versa describe the orthogonal light-clock based on the observations made by an observer in the same initial inertial frame as the light-clock. Therefore, the time interval between the sending and reception of the light signal is $\Delta t=2l/c$. The spacing of the time marks, i.e. the distance $\Delta x$ through which the screen in the accelerated frame falls in the time is $\Delta x=\Delta x_{n}=2lv/c$.
Analysis based on using standard measuring rods in an accelerated frame to measure its geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. We have ascertained the asymmetry of the meter stick’s length directly from the Langevin metric. This shows that the physical spacing of the time marks measured in the accelerated frame undergoes compress on in the direction of fixed stars motion. Because $\Delta x_{n}=\Delta x$ and $\Delta x=2vl/c$ has the physical meaning of the time rate indicated by a physical clock at rest in the inertial frame, this implies that the accelerated observer can observe the shortness of mark spacing $\Delta l_{n}=2lv\sqrt{1-v^{2}/c^{2}}]/c$. Here $\Delta l_{n}$ is the physical mark spacing measured in the accelerated frame. This thought experiment illustrates that there is no reciprocity in time dilation.
@@ 13.2 Invariance of the Two-Way Speed of Light and the Interference Phenomena
We must now discuss a certain feature of the phenomenon of interference. Let us consider a reflection and refraction in the context of special relativity. We are now in a position to understand what happens when light radiated by a moving emitter goes through a stationary material like glass. We see that what we have to do is to calculate the interference between the incident radiation and the radiation due to the material. At first glance, if one wants to calculate the refraction and the reflection in a single inertial frame one should take into account that the metrics, and consequently the electrodynamics equations, are different for a moving emitter and a stationary glass. So the possibility of calculating interference effects is not clear.
@@
Figure 48: Light radiated by a moving emitter reflects from a stationary mirror in the collinear geometry. There is interference between the incident radiation and the radiation due to the material.
@@
If the mirror is at rest and the emitter is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the same metric should be applied and kept in a consistent way for both incident and scattered waves. The requirement that the electrodynamics equations must be identical for both incident and reflected waves appears to be a paradox here. Indeed, we have two different metrics.
@@
We have already made a few remarks about two different metrics in the aberration of light problem. In the Section 4, we considered only the first-order approximation. We pointed out that the paradox is avoided because the peculiarity of the aberration of light geometry is that even after the Galilean transformation the electrodynamics equations in the absolute time coordinatization will have Maxwell’s form for the normally incident plane wave. Let us see what happens in the collinear geometry, Fig. 48.
@@
Here we summarize the general idea. We will show that the paradox is avoided because Minkowski metric and metric Eq. (11) give the same prediction. The point is that all methods to measure the interference, indeed, the standing wave (i.e. round-trip) measurements. In contrast, the deviation of the energy transport direction is a geometrical effect. Because we have empirical access only to the round-trip average speed of light the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast, the two-way speed of light, directly measurable along a round-trip, has physical meaning.
@@
We must now discuss a certain feature of the phenomenon of interference. It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different in the absolute time coordinatization. This is simply not true in physics. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase that varies with position in space. This affects the velocity and amplitude envelope of the single wave which results from the superposition of the two separate waves.
@@
The following simple analysis confirms these ideas. We will describe the effect by assuming that the metric is non-diagonal, Eq. 11. Let $\exp i(\omega t-kx)$ represent an incoming wave whose velocity is $\omega/k=c+v$. Similarly, let $\exp i(\omega t+k^{\prime}x+\phi)$ represent another out-coming (scattered radiation) wave of the same amplitude and the same frequency, a different velocity $\omega/k^{\prime}=c-v$ and different phase. The superposition of these two waves is represented by $\exp i(\omega t-kx)+\exp i(\omega t+k^{\prime}x+\phi)=2[\cos[(k+k^{\prime})x/2+\phi/2]]\exp i[\omega t-(k-k^{\prime})x/2+\phi/2]$. There is the cosine factor representing an amplitude envelope that is stationary in space and whose periodicity is inversely proportional to the difference in the propagating constants $k$ and $k^{\prime}$ of the two component waves. This can be written in a simpler form $2[\cos[\omega x/[c(1-v^{2}/c^{2})]+\phi/2]]\exp i[\omega t-xv\omega/[c^{2}(1-v^{2}/c^{2})]+\phi/2]$.
@@
Suppose that the source at rest is emitting waves at frequency $\omega_{0}$. In the lab frame after the Galilean transformation the velocity of the incoming wave is $c+v$. Thus if $\omega_{0}$ is the natural frequency, the observed in the lab frame frequency would be $\omega=\omega_{0}/[1-v/(c+v)]$. Therefore, the observed in the lab frame frequency is $\omega=\omega_{0}(1+v/c)$. The shift in frequency observed in the above situation is the well-known Doppler effect. Our equation for the superposition of two waves now looks like $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
@@
Suppose that an observer in the laboratory performs the standing wave measurement. We should examine what parts of the measured data depend on the choice of synchronization convention and what parts do not. We state that time oscillation has no intrinsic meaning - its meaning only being assigned by a convention. In particular, one can see the connection between the time shift $xv\omega_{0}(1+v/c)/c^{2}$ in $\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$ and the issue of distant clock synchrony. Note that the scale of time (frequency) is also unrecognizable from a physical viewpoint.
@@
Suppose that the laboratory observer measures the wavelength of the standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ gives the connection between the spatial line element and the metric. With the help of Eq.(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the case of the Langevin metric. This is explained as due to Lorentz contraction of the measuring rods. Then, when the measured data is analyzed, the laboratory observer finds that the wave number is equal to $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/\sqrt{1-v/c}$. The observer finds that the wavelength of radiation from the moving source (the source moves towards the observer) is decreased by the factor $\sqrt{1-v/c}/\sqrt{1+v/c}$. We see that it is the same factor that we can obtain by assuming that the metric is diagonal. In other words, the laboratory observer will measure the same two-way speed of light, irrespective of the choice of metric.
@@
The reason why the laboratory observer will measure the same wavelength is that the method is based on the phenomenon of interference. There is a fact governing all interference phenomena. The law is called the relativistic invariance of phase. It states that there is a certain quantity, which we call phase, that does not change under any reference frame transformations. The relativistic invariance of phase is enormously useful in making an analysis (as we will illustrate in a few examples in the next section) without knowing all the formulas.